A solution is prepared by adding 1.59 mg of an unknown protein to enough water to make a 7.3 mL solution. The osmotic pressure of the solution is measured to be 1E-3 bar at 25 C. What is the calculated molar mass of the protein?
1
Expert's answer
2019-09-17T04:54:36-0400

O.P. = iMRT

O.P. = osmotic pressure = 1x10-3 bar

i = van't Hoff factor = 1 for a protein

M = molarity = ?

T = temp = 25 + 273 = 298K

R = gas constant = 0.08314 Lbar/Kmol

Solving for M, we have...

M = O.P./(i)(R)(T) = (1x10-3 bar)/(1)(0.08314 Latm/Kmol)(298K)

M = 4.0x10-8 moles/L

moles = 4.0x10-8 mole/L x 0.0073 L = 5.5x10-6 moles

MW = 1.59 mg x 1 g/1000 mg/5.5x10-6 moles = 289 g/mole